OpenStudy (18jonea):

Two vehicles are approaching an intersection perpendicular to each other. One vehicle has a mass of 1,125 kg and is traveling at 75.0 km/h. The other vehicle has a mass of 1,575 kg and is traveling at 60.0 km/h. What is the magnitude of the momentum after an inelastic collision? @mathmate

OpenStudy (mathmate):

Hints: 1. For inelastic collisions, m1u1+m2u2=(m1+m2)v u1, u2=velicities before impact v=common velocity after inelastic collision. 2. Momentum is a vector quantity, so calculate momentum in each orthogonal (perpendicular) direction and add the two components vectorially to find the magnitude.